### v1

# BeeManc at the PLABA Track of TAC-2024: RoBERTa for task 1 LLaMA3.1 and GPT-4o for task 2

Abstract
This report is the system description of the BeeManc team for shared task Plain Language Adaptation of Biomedical Abstracts (PLABA) 2024. This report contains two sections corresponding to the two sub-tasks in PLABA 2024. In task one, we applied fine-tuned ReBERTa-Base models to identify and classify the difficult terms, jargon and acronyms in the biomedical abstracts and reported the F1 score. Due to time constraints, we didn’t finish the replacement task. In task two, we leveraged Llamma3.1-70B-Instruct and GPT-4o with the one-shot prompts to complete the abstract adaptation and reported the scores in BLEU, SARI, BERTScore, LENS, and SALSA. From the official Evaluation from PLABA-2024 on Task 1A and 1B, our much smaller fine-tuned RoBERTa-Base model ranked 3rd and 2nd respectively on the two sub-task, and the 1st on averaged F1 scores across the two tasks from 9 evaluated systems. Our share our fine-tuned models and related resources at https://github.com/HECTA-UoM/PLABA2024
1 Background
Health literacy, or the ability of individuals to comprehend and apply health information for informed decision-making, is one of the central focuses of the Healthy People 2030 framework in the US. Even though biomedical information is highly accessible online, patients and caregivers often struggle with language barriers, even when the content is presented in their native language.
The shared task PLABA aims to harness advances in deep learning to empower the automatic simplification of complex scientific texts into language that is more understandable for patients and caregivers. Despite substantial obstacles to effective implementation, the goal of the PLABA track is to improve health literacy by translating biomedical abstracts into plain language, making them more accessible and understandable to the general public  [^1] https://bionlp.nlm.nih.gov/plaba2024/. Following our participation on the PLABA 2023 shared task using large language models (LLMs) such as ChatGPT, BioGPT, and Flan-T5, and Control Mechanisms Li et al. (2024), in this work, we introduce our system participation to the PLABA 2024. Instead of end-to-end biomedical abstract simplification as in PLABA-2023, in this year, PLABA-2024 introduced more granular-steps, including Term Replacement for Task-1 and Complete Abstract Adaption for Task-2, which we will describe in detail for our methodologies via fine-tuning RoBERTa-Base model for Task-1 and prompting LLMs (Llama and ChatGPT) for Task-2.
2 PLABA 2024 Task 1: Term Replacement
2.1 Introduction for task 1
Task 1 does not require a full adaptation process. Instead, the system will identify challenging terminology, determine appropriate strategies for addressing these terms, and offer suitable replacements.
This task is split into three subtasks:
1A: Identifying non-consumer terms
This is a Name Entity Recognition (NER) task. The objective is to to extract a list of exact phrases from the given abstracts, each representing a concept that a consumer would not understand.
1B: Classifying replacement
This is a multi-class multi-label token classification task. Systems should identify any and all types of replacements that would be appropriate for each identified non-consumer term.
1C: Generation
Provide text for each positive label from 1B (except for the ”omitted” label).
2.2 Methodology
Figure 1: System overview for task 1A and 1B
2.2.1 Corpus
In this task, we leveraged the PLABA corpus Attal et al. (2022), which contains biomedical queries and the abstracts of top-ten papers corresponding to the query from PubMed. The sentences in each abstract are manually aligned and simplified by experts. In addition, the difficult phrases are tagged with proper labels indicating the replacement method. The original dataset is in JSON format and one sample is shown in Fig2.
There are five different kinds of replacement labels, and the task requires systems to assign these labels non-exclusively to the phrases.
  • Substituted: the term is jargon with a common alternative (e.g. ”myocardial infarction” can be ”heart attack”).
  • Explained: there is no alternative, or the term is important to the topic, and it should be explained. For example: ”This study looked at treatments for sleep apnoea (when you stop breathing while sleeping).”
  • Generalized: the term can be replaced with a more general category without losing its significance. For example: ”Clearing of the infection is confirmed with a [Nucleic Acid Amplification Test (NAAT)] common lab test.”
  • Exemplified: the term has a specific example that would give a general audience an idea of what it is. For example: ”Depression is common in people with neurodegenerative diseases (like Parkinson’s).”
  • Omitted: The term is not relevant to understanding the sentence or too technical to explain, and does not need to appear in a consumer version.
To adapt this data to a form that is suitable for an NER task, we need to map these labels to the phrases in the sentences.
2.2.2 Data Preprocessing
We introduce how we modify the training data to a format for our task setting. For each sentence, terms that are difficult to read are given human-annotated replacement labels. Terms can be a single word or a phrase, and label types can also be multiple.
For task 1A, we define it as a word-level binary NER task. Because this subtask is covered by subtask 1B, we did not develop a system to specifically address this task. The output for task 1A is the words that are assigned not the O label in task 1B.
For Task 1B, labels must be mapped to subword-level labels during model fine-tuning, as the model’s tokenizer segments sentences into subword units. The original training data shown in Figure 2 contains phrase-level annotations, which are not directly compatible with the subword-level classification required by our NER model. To address this issue, we implemented a two-step label transformation process. First, we decomposed the annotated phrases into individual words, applying the BIO tagging scheme: the first word of each phrase receives a B-[tag] (Beginning), while subsequent words receive I-[tag] (Inside) labels. Second, to accommodate the model’s tokenizer, which further segments words into subwords, we developed the following labelling strategy:
  • For words originally labeled with B-[tag]: the first subword maintains the B-[tag], while subsequent subwords receive I-[tag] labels
  • For words originally labeled with I-[tag]: all resulting subwords inherit the I-[tag] label
  • Context words outside the annotated phrases are labelled with O (Outside)
This approach ensures consistent label propagation from phrase-level annotations to subword-level classifications needed to train the NER model.
The classification uses the following label types:
  • SUBSTITUTE
  • EXPLAIN
  • EXEMPLIFY
  • GENERALIZE
  • OMIT
Each type follows the BIO tagging scheme with Beginning (B-) and Inside (I-) variants, plus an additional Outside (O) label, yielding eleven labels in total. Finally, we split the data into training, validation, and test sets with a ratio of 8:1:1, resulting in 836 sentences for training, 105 for validation, and 105 for testing. Table 1 shows the distribution of the label for each set.
Figure 2: an Example of the corpus
Label Train Valid Test
O 17703 2005 2411
B-SUBSTITUTE 1989 264 239
I-SUBSTITUTE 1385 181 191
B-EXPLAIN 1047 160 122
I-EXPLAIN 549 68 62
B-GENERALIZE 317 50 39
I-GENERALIZE 419 86 52
B-EXEMPLIFY 44 6 6
I-EXEMPLIFY 37 2 7
B-OMIT 309 45 50
I-OMIT 314 38 79
Table 1: Label Counts for Training, Validation, and Test Sets
2.2.3 Model Description
For this multi-label, multi-class classification task, we selected Robustly Optimized BERT Pretraining Approach (RoBERTa) Liu et al. (2019) as our model of choice. RoBERTa is a transformer-based language model. It enhances the pretraining methodology of Bidirectional Encoder Representations from Transformers (BERT) to achieve superior performance in natural language processing tasks. Specifically, RoBERTa removes the next-sentence prediction objective, increases the amount of training data, and employs larger batch sizes and learning rates during training. These optimizations enable RoBERTa to outperform BERT on several benchmarks, demonstrating its advanced capabilities in understanding and generating human language. We used the open-source code from INSIGHTBuddy-AI for RoBERTa fine-tuning Romero et al. (2024).
2.3 Experiments
Figure 3: The confusion matrices for the results on the test set. The figure shows the confusion matrix for each of the labels.
Figure 4: The confusion matrix between all labels on the test set.
Fine tuning process
We add a linear layer after RoBERTa’s final hidden layer. The system converts the output for each token into a set of probabilities for the 11 classes using the sigmoid function. We use Binary Cross Entropy as our loss function, specifically implementing it through PyTorch’s BCEWithLogitsLoss [^2] https://pytorch.org/docs/stable/generated/torch.nn.BCEWithLogitsLoss.html#bcewithlogitsloss. A threshold of 0.5 is applied after the sigmoid function to determine if a token belongs to a particular label. Because we use sigmoid rather than softmax, each token can be assigned multiple labels. The detailed fine-tuning settings are shown in Table 5.
Post-processing
After the system assigns labels to subwords, we merge them back into word-level labels. Our strategy for combining subwords into words follows these rules:
  1. 1.
    A subword token beginning with the special character ”Ġ” indicates the start of a new word.
  2. 2.
    If the subsequent token also begins with ”Ġ”, the current token is treated as a complete word and retains its label.
  3. 3.
    When a token with a leading ”Ġ” is followed by one or more tokens without ”Ġ”, these tokens are merged into a single word. The merged word inherits the label from the initial ”Ġ” token. For example, the tokens [’Ġdoc’, ’us’, ’ate’] with labels [’B-SUBSTITUTE’, ’B-EXPLAIN’], [’I-SUBSTITUTE’, ’I-EXPLAIN’], [’I-SUBSTITUTE’, ’I-EXPLAIN’] merge to form ’docusate’ with the labels [’B-SUBSTITUTE’, ’B-EXPLAIN’].
  4. 4.
    If a word’s first token is labelled ’O’ but its later subwords have other labels, the whole word takes the first non-’O’ label. For example: [’Ġdoc’(’O’), ’us’(’O’), ’ate’(’I-SUBSTITUTE’)] becomes ’docusate’(’I-SUBSTITUTE’).
  5. 5.
    Special sequence tokens (<s >, </s >) added by RoBERTa’s tokenizer are removed as they are not part of the actual text content.
As the output of the format is phrase level, we need to merge the words back into phrases, here is our merging strategy for forming phrases:
  1. 1.
    When a token has a B-[tag], we combine it with all following I-[tag] tokens to form a phrase. For example: [’repeat’(B-SUBSTITUTE), ’infect’(I-SUBSTITUTE), ’ion’(I-SUBSTITUTE), ’rates’(I-SUBSTITUTE)] forms the phrase ”repeat infection rates” with label SUBSTITUTE.
  2. 2.
    If we find an I-[tag] without a preceding B-[tag], we trace back to the first token starting with ”Ġ” to mark the phrase beginning.
The output consists of identified difficult phrases and their corresponding replacement labels at the phrase level.
2.4 Results/Discussion
We report Automatic Evaluation of the word level evaluation for task 1B in table 2.
F1-micro Precision-micro Recall-micro
word level(w/ [’O’] label) 0.836 0.858 0.819
word level(w/o [’O’] label) 0.413 0.546 0.417
Table 2: System performance on the test set for task 1B.
The RoBERTa-base model’s performance shows significant limitations. While the word-level NER achieves a micro-F1 score of 0.836 when including [’O’] labels, this score drops substantially to 0.413 when excluding them. The confusion matrices for all 11 labels are shown in Fig 3. This performance degradation primarily stems from the model’s poor prediction of low-frequency labels in the training data, particularly EXEMPLIFY, OMIT, and GENERALIZE. The limited training instances for these labels appear insufficient for effective model fine-tuning. Analysis of the training process reveals that the model rarely correctly classified these minority labels, suggesting room for improvement in handling class imbalance.
3 PLABA 2024 Task 2: Complete Abstract Adaptation
3.1 Introduction
Task 2 involves the end-to-end adaptation of biomedical abstracts for a general audience by utilizing plain language. Given a collection of abstracts as the source material, the system will generate a corresponding plain language adaptation as the output for each sentence within the source text.
3.2 Methodology
To generate simplified versions of abstract sentences, we employed the Llama-3.1-70B Dubey et al. (2024) instruction tuned version  [^3] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct and GPT-4o OpenAI et al. (2024) with one-shot prompt only.
3.2.1 Corpus
In this shared task, we stick with the same PLABA corpus Attal et al. (2022), which is introduced in section 2.2.1, but it is mostly used for evaluation only. In the corpus, there are 40 abstracts containing 7604 sentences in total with a reference or references.
3.2.2 Metrics
In this shared task, we mainly rely on automatic evaluation metrics, which can be classified into reference-based and reference-less metrics. In reference-based metrics, BLEU Papineni et al. (2002), SARI xu- (2016), BERTScore Zhang* et al. (2020) and LENS Maddela et al. (2023) are chosen due to their popularity in text simplification areas. SALSA Heineman et al. (2023), which is a new-developed metric, is used for reference-less evaluation.
BLEU
The Bilingual Evaluation Understudy (BLEU) score Papineni et al. (2002) is a metric widely used in machine translation to evaluate the quality of the generated text. As text simplification can be regarded as monolingual text translation, it is also accepted and used in the early stages of text simplification studies. It measures the n-gram (from 1 to 4) text similarity between the generated text and the single or multiple references.
SARI
The System Output Against Reference and Input (SARI) score xu- (2016) is a metric designed to evaluate text simplification tasks. Similar to BLEU, SARI also measures the n-gram similarity between the output and references. Additionally, the score is determined based on three operations: adding, deleting, and retaining words from the original sentence. Each operation is assessed by comparing it to one or more reference simplified texts to measure how closely the simplified text matches expert simplifications. As a result, it better aligns with human evaluations in text simplification. simplification.
BERTScore
Unlike the two above-mentioned metrics, BERTSCore Zhang* et al. (2020) measures the semantic similarity between the output and reference instead of n-gram similarity. By leveraging the deep contextualized embeddings from BERT Devlin et al. (2019), it can measure the semantic similarity among the tokens. Then it calculates the sentence similarity by token matching and maximising the matrix between output and references.
LENS
Learned Evaluation of Non-Canonical Sequences (LENS) Maddela et al. (2023) is yet another metric for evaluating text generation tasks. The embedded LENS machine learning model in LENS makes it different from the previous metrics. It evaluates the generated text based on semantic and contextual similarity, which improves the adaptability and accuracy in tasks where the generated text may differ in surface form from reference text but still be semantically correct.
SALSA
Success and FAilure driven Linguistic Simplification Annotation (SALSA) Heineman et al. (2023) is an edit-based human annotation framework. By training LENS Maddela et al. (2023) on 19 thousand edit annotations gathered through SALSA, the authors developed LENS-SALSA, which we refer to simply as SALSA in this report. The main difference between SALSA and LENS is that SALSA can measure performance based on self-predicted references, which is beneficial for evaluating the test set.
3.3 Experiment
3.3.1 prompting process
To generate simplified sentences in the abstract, we started with a simple 1-shot prompt that has one example in the prompt. Then based on the output of test cases, we kept asking the model to self-adjust the prompt until it almost generated the expected outcome. We list the prompts we used in Tables LABEL:tab:promptgpt4o and LABEL:tab:promptllama in the appendix.
3.3.2 post-processing
Sometimes models will output some pre-words such as ”Here’s a simplified version: ” and ”**Simplified**:”, we checked the outputs and used the rule to delete these pre-words. Some outputs have detailed explanations for how and why did I simplify this sentence. For this kind of output, we used regular expressions to extract the simplified sentence as possible.
3.4 Results/Discussion
Automatic Evaluation
Model SALSA LENS SARI BERT(F1)
Llama 3.1 instruct-70B 73.89 52.79 36.37 0.88
GPT-4o 74.28 64.75 37.72 0.92
Table 3: System performance on the evaluation set across SALSA, LENS, SARI and BERT for BERTScore.
Model SALSA
[…]
Table 4: System performance on the test set under the SALSA (95% confidence interval).
Tables 3 and 4 show the performance difference between the two models across the metrics. In Table 3, all four metrics consistently favour GPT-4o over Llama and shows some credibility of SALSA as a reference-less metric. Similarly in Table 4, the two models show an aligned performance gap, compared to Table 3.
4 Official Ranking from PLABA2024
Figure 5 shows the system ranking from PLABA2024 official organisers on Task 1A and 1B using F1 scores, 1C being not popular among the submissions. There are overall 9 systems evaluated by PLABA 2024 Ondov et al. (2024). Out of the 9 systems, our team UM ranks the third for Task 1A, and the 2nd for Task 1B via F1 scores. There are a few intersting findings from the results.
  • Firstly, it is worth to note that for both tasks, our model scores are very close to the highest system, e.g. for Task 1A, our F1 score is 0.4787 using Roberta-Base, after the first two systems 0.5036 from ReBERTa-GBC and 0.4885 from Gemini-1.5. For Task 1B, our F1 score 0.7765 is even closer to the top-1 score 0.7788 from MLPClassifier system.
  • Secondly, our average F1 score across Task 1A and 1B is the highest among all 9 systems, 0.6276=(0.4787+0.7765)/2.
  • Thirdly, It is interesting to see that much smaller models can outperform the extra-large LLMs on Task 1A and 1B, e.g. our RoBERTa-base fine-tuned wins Mistral, GPT, Gemini-1.5 on these two tasks. This is related to the findings from Li et al. (2024) for PLABA 2023 shared task.
Figure 5: System Ranking from PLABA2024 submissions: Task 1A and 1B Ondov et al. (2024)
5 Conclusion
In this report, we present the two systems submitted by the BeeManc team for two different tasks in PLABA 2024. For Task 1, we submitted a classification system based on Roberta for Task 1a and 1b. The system’s performance at the word level was not outstanding, highlighting the limitations of simple classification models in multi-label, multi-class tasks. For Task 2, we submitted end-to-end generation results based on LLaMa 3.1 and GPT-4o for the generation task. On the Evaluation set, GPT-4o performed slightly better than LLaMa. In the unsupervised evaluation on the test set, GPT-4o also outperformed LLaMa 3.1. Our findings demonstrate the potential of open-source models in text simplification tasks within the medical domain.
6 Acknowledgment
The authors wrote the initial draft, used Generative AI to revise/polish it, then the authors reviewed the generated content. The authors used ChatGPT to refine the paper writing.

### v2

@@
-BeeManc at the PLABA Track of TAC-2024: RoBERTa for task 1 LLaMA3.1 and GPT-4o for task 2
+BeeManc at the PLABA Track of TAC-2024: RoBERTa for Task 1 – LLaMA3.1 and GPT-4o for Task 2
 Abstract
 This report is the system description of the BeeManc team for shared task Plain Language Adaptation of Biomedical Abstracts (PLABA) 2024.
-This report contains two sections corresponding to the two sub-tasks in PLABA 2024.
+This report contains two sections corresponding to the two sub-tasks in PLABA-2024.
 In task one, we applied fine-tuned ReBERTa-Base models to identify and classify the difficult terms, jargon and acronyms in the biomedical abstracts and reported the F1 score.
@@
 In task two, we leveraged Llamma3.1-70B-Instruct and GPT-4o with the one-shot prompts to complete the abstract adaptation and reported the scores in BLEU, SARI, BERTScore, LENS, and SALSA.
-From the official Evaluation from PLABA-2024 on Task 1A and 1B, our much smaller fine-tuned RoBERTa-Base model ranked 3rd and 2nd respectively on the two sub-task, and the 1st on averaged F1 scores across the two tasks from 9 evaluated systems.
-Our share our fine-tuned models and related resources at https://github.com/HECTA-UoM/PLABA2024
+From the official Evaluation from PLABA-2024 on Task 1A and 1B, our much smaller fine-tuned RoBERTa-Base model ranked 3rd and 2nd respectively on the two sub-tasks, and the 1st on averaged F1 scores across the two tasks from 9 evaluated systems.
+Our LLaMA-3.1-70B-instructed model achieved the highest Completeness score for Task-2.
+We share our source codes, fine-tuned models, and related resources at https://github.com/HECTA-UoM/PLABA-2024
 1 Background
@@
 The shared task PLABA aims to harness advances in deep learning to empower the automatic simplification of complex scientific texts into language that is more understandable for patients and caregivers.
-Despite substantial obstacles to effective implementation, the goal of the PLABA track is to improve health literacy by translating biomedical abstracts into plain language, making them more accessible and understandable to the general public  [^1] https://bionlp.nlm.nih.gov/plaba2024/.
-Following our participation on the PLABA 2023 shared task using large language models (LLMs) such as ChatGPT, BioGPT, and Flan-T5, and Control Mechanisms Li et al.
-(2024), in this work, we introduce our system participation to the PLABA 2024.
-Instead of end-to-end biomedical abstract simplification as in PLABA-2023, in this year, PLABA-2024 introduced more granular-steps, including Term Replacement for Task-1 and Complete Abstract Adaption for Task-2, which we will describe in detail for our methodologies via fine-tuning RoBERTa-Base model for Task-1 and prompting LLMs (Llama and ChatGPT) for Task-2.
-2 PLABA 2024 Task 1: Term Replacement
+Despite substantial obstacles to effective implementation, the goal of the PLABA track is to improve health literacy by translating biomedical abstracts into plain language, making them more accessible and understandable to the general public  [^1] https://bionlp.nlm.nih.gov/PLABA-2024/.
+Following our participation on the PLABA-2023 shared task using large language models (LLMs) such as ChatGPT, BioGPT, and Flan-T5, and Control Mechanisms Li et al.
+(2024), in this work, we introduce our system participation to the PLABA-2024.
+Instead of end-to-end biomedical abstract simplification as in PLABA-2023, in this year, PLABA-2024 introduced more granular-steps, including Term Replacement for Task-1 and Complete Abstract Adaption for Task-2, which we will describe in detail for our methodologies via fine-tuning RoBERTa-Base model for Task-1 and prompting LLMs (LLaMa-3.1-70B and GPT4o) for Task-2.
+2 PLABA-2024 Task 1: Term Replacement
 2.1 Introduction for Task 1
@@
 1C: Generation
-Provide text for each positive label from 1B (except for the ”omitted” label).
+Provide text for each positive label from 1B (except for the “OMIT” label). For this task, there aren’t popular submissions, so we will skip this task.
 2.2 Methodology
@@
 Each type follows the BIO tagging scheme with Beginning (B-) and Inside (I-) variants, plus an additional Outside (O) label, yielding eleven labels in total.
-Finally, we split the data into training, validation, and test sets with a ratio of 8:1:1, resulting in 836 sentences for training, 105 for validation, and 105 for testing.
+Finally, we split the official PLABA-2024 data into training, validation, and test sets with a ratio of 8:1:1, resulting in 836 sentences for training, 105 for validation, and 105 for testing.
 Table 1 shows the distribution of the label for each set.
-Figure 2: an Example of the corpus
+Figure 2: Example of training corpus for Task 1
 Label Train Valid Test
@@
 I-OMIT 314 38 79
-Table 1: Label Counts for Training, Validation, and Test Sets
+Table 1: Label Counts for Training, Validation, and Test Sets for Task 1B: 11 labels
 2.2.3 Model Description
-For this multi-label, multi-class classification task, we selected Robustly Optimized BERT Pretraining Approach (RoBERTa) Liu et al.
+For this multi-label, multi-class classification task, we selected the Robustly Optimized BERT Pretraining Approach (RoBERTa) Liu et al.
 (2019) as our model of choice.
@@
 2.3 Experiments
-Figure 3: The confusion matrices for the results on the test set. The figure shows the confusion matrix for each of the labels.
-Figure 4: The confusion matrix between all labels on the test set.
+Figure 3: The confusion matrices for the results on the test set of Task-1B for each label category.
+Figure 4: The confusion matrix between all labels on the test set (our 8:1:1 split) for Task 1B.
 Fine tuning process
@@
 Post-processing
-After the system assigns labels to subwords, we merge them back into word-level labels. Our strategy for combining subwords into words follows these rules:
+After the system assigns subword labels, we merge them into word-level labels. Our strategy for combining subwords into words follows these rules:
   1. 1.
-    A subword token beginning with the special character ”Ġ” indicates the start of a new word.
+    A subword token beginning with the special character “Ġ” indicates the start of a new word.333“Ġ” is the special token from RoBERTa tokenizer; it is automatically attached to a subword when we do “tokenizer.decode” function.
   2. 2.
-    If the subsequent token also begins with ”Ġ”, the current token is treated as a complete word and retains its label.
+    If the subsequent token also begins with “Ġ”, the current token is treated as a complete word and retains its label.
   3. 3.
-    When a token with a leading ”Ġ” is followed by one or more tokens without ”Ġ”, these tokens are merged into a single word.
-The merged word inherits the label from the initial ”Ġ” token.
-For example, the tokens [’Ġdoc’, ’us’, ’ate’] with labels [’B-SUBSTITUTE’, ’B-EXPLAIN’], [’I-SUBSTITUTE’, ’I-EXPLAIN’], [’I-SUBSTITUTE’, ’I-EXPLAIN’] merge to form ’docusate’ with the labels [’B-SUBSTITUTE’, ’B-EXPLAIN’].
+    When a token with a leading “Ġ” is followed by one or more tokens without “Ġ”, these tokens are merged into a single word.
+The merged word inherits the label from the initial “Ġ” token.
+For example, the tokens [‘Ġdoc’, ‘us’, ‘ate’] with labels [’B-SUBSTITUTE’, ’B-EXPLAIN’], [’I-SUBSTITUTE’, ’I-EXPLAIN’], [’I-SUBSTITUTE’, ’I-EXPLAIN’] merge to form ’docusate’ with the labels [’B-SUBSTITUTE’, ’B-EXPLAIN’].
   4. 4.
@@
 2.4 Results/Discussion
-We report Automatic Evaluation of the word level evaluation for task 1B in table 2.
+We report Automatic Evaluation at the word level for task 1B in table 2.
 F1-micro Precision-micro Recall-micro
@@
 word level(w/o [’O’] label) 0.413 0.546 0.417
-Table 2: System performance on the test set for task 1B.
+F1-macro Precision-macro Recall-macro
+word level(w/ [’O’] label) 0.274 0.276 0.276
+word level(w/o [’O’] label) 0.206 0.211 0.208
+Table 2: System performance on the test set (our 8:1:1 split of Attal et al. (2023)) for task 1B.
 The RoBERTa-base model’s performance shows significant limitations.
@@
 Analysis of the training process reveals that the model rarely correctly classified these minority labels, suggesting room for improvement in handling class imbalance.
-3 PLABA 2024 Task 2: Complete Abstract Adaptation
+The data and label distribution in Table 1 also reveals that the lower macro score is due to the zero accuracy on the low-ratio labels (Figure 4 lower part of the diagonal line).
+Even though these labels have a small appearance in the entire dataset, macro score assigns each label class the same weight, bringing down the overall macro scores.
+However, the overall performance of the system is still reasonable instead of nonsense, which can be seen in the upper part of Table 2, where the micro F1 score without the O label is still 0.413 on all special labels.
+3 PLABA-2024 Task 2: Complete Abstract Adaptation
 3.1 Introduction
@@
 3.2 Methodology
-To generate simplified versions of abstract sentences, we employed the Llama-3.1-70B Dubey et al.
-(2024) instruction tuned version  [^3] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct and GPT-4o OpenAI et al.
-(2024) with one-shot prompt only.
+To generate simplified versions of abstract sentences, we employed the Llama-3.1-70B dub (2024) instruction tuned version  [^4] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct and GPT-4o ope (2024) with one-shot prompt only.
 3.2.1 Corpus
@@
 In reference-based metrics, BLEU Papineni et al.
-(2002), SARI xu- (2016), BERTScore Zhang* et al.
+(2002), SARI Xu et al.
+(2016), BERTScore Zhang* et al.
 (2020) and LENS Maddela et al.
@@
 SARI
-The System Output Against Reference and Input (SARI) score xu- (2016) is a metric designed to evaluate text simplification tasks.
+The System Output Against Reference and Input (SARI) score Xu et al.
+(2016) is a metric designed to evaluate text simplification tasks.
 Similar to BLEU, SARI also measures the n-gram similarity between the output and references.
@@
 For this kind of output, we used regular expressions to extract the simplified sentence as possible.
-3.4 Results/Discussion
+3.4 Results and Discussion
 Automatic Evaluation
@@
 GPT-4o 74.28 64.75 37.72 0.92
-Table 3: System performance on the evaluation set across SALSA, LENS, SARI and BERT for BERTScore.
+Table 3: System performance on the full training set across SALSA, LENS, SARI and BERT for BERTScore. The training set is provided by PLABA dataset Attal et al. (2023)
 Model SALSA
@@
 […]
-Table 4: System performance on the test set under the SALSA (95% confidence interval).
+Table 4: System performance on the test set under the SALSA (95% confidence interval). The test set is provided by PLABA-2024, and is different from the training set.
 Tables 3 and 4 show the performance difference between the two models across the metrics.
-In Table 3, all four metrics consistently favour GPT-4o over Llama and shows some credibility of SALSA as a reference-less metric.
+In Table 3, all four metrics consistently favour GPT-4o over Llama and show some credibility of SALSA as a reference-less metric.
 Similarly in Table 4, the two models show an aligned performance gap, compared to Table 3.
-4 Official Ranking from PLABA2024
-Figure 5 shows the system ranking from PLABA2024 official organisers on Task 1A and 1B using F1 scores, 1C being not popular among the submissions.
-There are overall 9 systems evaluated by PLABA 2024 Ondov et al.
+Comparing these two tables (3 and 4):
+  • With (source, output, reference) set, the difference of LLaMa-3.1-70B and GPT-4o on different metrics is not that large margin, except for LENS, as in Table 3.
+  • With (source, output) no reference: the SALSA score gap increased dramatically, as in Table 4.
+4 Official Ranking from PLABA-2024
+4.1 Task-1 (A and B)
+Table 5 shows the system ranking from PLABA-2024 official organisers on Task 1A and 1B using F1 scores, 1C being not popular among the submissions.
+There are overall 9 systems evaluated by PLABA-2024 Ondov et al.
 (2024).
@@
 For Task 1B, our F1 score 0.7765 is even closer to the top-1 score 0.7788 from MLPClassifier system.
-  • Secondly, our average F1 score across Task 1A and 1B is the highest among all 9 systems, 0.6276=(0.4787+0.7765)/2.
+In general, our fine-tuned system is almost as good as the top-1 system at identifying the complicated term, but we are much better at classifying the next to-do step in the term processing (11 labels = 1 + 5x2) compared to the top-1 system for Task-1A (0.7765 vs 0.6838) winning almost 10 absolute point difference.
+It is also interesting to see that the Top-1 system for Task-1B (F1 0.7788) from MLPClassifier, produced the lowest F1 score for Task-1A, i.e.
+0.0459, which means it actually cannot identify most of the complicated terms (only made into 5%).
+  • Secondly, our averaged F1 score across Task 1A and 1B is the highest among all 9 systems, 0.6276=(0.4787+0.7765)/2. It is also the only system whose Averaged F1 score is above 0.60.
   • Thirdly, It is interesting to see that much smaller models can outperform the extra-large LLMs on Task 1A and 1B, e.g.
@@
 This is related to the findings from Li et al.
-(2024) for PLABA 2023 shared task.
-Figure 5: System Ranking from PLABA2024 submissions: Task 1A and 1B Ondov et al. (2024)
-5 Conclusion
-In this report, we present the two systems submitted by the BeeManc team for two different tasks in PLABA 2024.
+(2024) for the PLABA-2023 shared task.
+Team Run Rank Task 1A (F1) Task 1B (F1) Average Score BU MLPClassifier-identify-classify-replace-v1 1 0.0459 0.7788 0.4124 CLAC mistral 1 0.4410 0.6663 0.5537 CLAC gpt 2 0.3767 0.3795 0.3781 IITH First 1 0.1956 0.7014 0.4485 UM Roberta-base 1 0.4787 0.7765 0.6276 ntu_nlp gemini-1.5-pro_demon5_replace-demon5 1 0.4885 0.6335 0.5610 ntu_nlp gemini-1.5-flash_demon5_replace-demon5 2 0.4431 0.6544 0.5488 ntu_nlp gpt-4o-mini_demon5_replace-demon5 3 0.4518 0.6197 0.5358 Yseop roberta-gbc 1 0.5036 0.6838 0.5937
+Table 5: Task 1A and Task 1B (F1) Results: Roberta-base highest Averaged-ScoreOndov et al. (2024)
+4.2 Task-2
+For Task 2, even though GPT4o outscored LLaMa-3.1-70B in our model development phase, to understand better how open-science LLMs perform on biomedical text simplification tasks, we selected LLaMa-3.1-70B as our prioritised output over the GPT4o-instruct for official human evaluation in the PLABA-2024.
+As shown in Table 6, it produced interesting outcomes with the highest Completeness score (0.9481) vs the lowest Accuracy score (0.6088).
+According to the human evaluation guidelines from PLABA-2024, Accuracy is “Outputs should contain the accurate information”, while Completeness is “Outputs should seek to minimize information lost from the original text”.
+So, LLaMa-3.1-70B-instructed can do its best to avoid information loss among the 8 evaluated systems, but it did not produce the most accurate outputs.
+These are somehow contradicting findings but it might be because the model produced some misleading outputs in addition to keeping the source information.
+We need further qualitative analysis on this assumption.
+Run Accuracy Completeness Simplicity Brevity Final avg.
+LLaMA-8B-4bit-MedicalAbstract-seq-to-seq-v1 0.8545 0.9180 0.6533 0.3978 0.7059 LLaMa 3.1 70B instruction 2nd run 0.6088 0.9481 0.5561 0.5004 0.6533 TREC2024_SIB_run3 0.7506 0.7716 0.7997 0.6484 0.7426 UAms-ConBART-Cochrane 0.9534 0.9398 0.6851 0.6171 0.7988 gpt35_dspy 0.9167 0.9386 0.6974 0.5478 0.7751 mistral-FINAL 0.6694 0.7552 0.5149 0.2875 0.5567 plaba_um_fhs_sub1 0.9067 0.9138 0.7814 0.6799 0.8205 task2_moa_tier1_post 0.8982 0.9444 0.7246 0.5284 0.7739
+Table 6: PLABA-2024 Task 2 results: LLaMa-3.1 with the highest Completeness Ondov et al. (2024)
+5 Conclusion and Future Work
+In this report, we present the two systems submitted by the BeeManc team for two different tasks in PLABA-2024.
 For Task 1, we submitted a classification system based on Roberta for Task 1a and 1b.
@@
 Our findings demonstrate the potential of open-source models in text simplification tasks within the medical domain.
+Regarding LLaMa-3.1-70B-instructed model producing the highest Completeness score but the lowest Accuracy according to the official human evaluation, we plan to carry out further qualitative analysis on its output to see if the system output contains misleading or hallucinated information in addition to the secured source knowledge.
+We can also use some specific human-in-the-loop metrics such as HOPE Gladkoff and Han (2022) to categorise the error types.
 6 Acknowledgment
-The authors wrote the initial draft, used Generative AI to revise/polish it, then the authors reviewed the generated content. The authors used ChatGPT to refine the paper writing.
+We thank Nicolo and Sam for their valuable discussion and feedback during this project with their expertise from the experience of PLABA-2023.
+The authors wrote the initial draft, used Generative AI to revise/polish it, then the authors reviewed the generated content.
+The authors used ChatGPT to refine the paper writing.

### v4

@@
-BeeManc at the PLABA Track of TAC-2024: RoBERTa for Task 1 – LLaMA3.1 and GPT-4o for Task 2
+MaLei at the PLABA Track of TREC 2024: RoBERTa for Term Replacement – LLaMA3.1 and GPT-4o for Complete Abstract Adaptation
 Abstract
-This report is the system description of the BeeManc team for shared task Plain Language Adaptation of Biomedical Abstracts (PLABA) 2024.
+This report is the system description of the MaLei team (Manchester and Leiden) for the shared task Plain Language Adaptation of Biomedical Abstracts (PLABA) 2024 (we had an earlier name BeeManc following last year), affiliated with TREC2024 (33rd Text REtrieval Conference https://ir.nist.gov/evalbase/conf/trec-2024).
 This report contains two sections corresponding to the two sub-tasks in PLABA-2024.
-In task one, we applied fine-tuned ReBERTa-Base models to identify and classify the difficult terms, jargon and acronyms in the biomedical abstracts and reported the F1 score.
-Due to time constraints, we didn’t finish the replacement task.
-In task two, we leveraged Llamma3.1-70B-Instruct and GPT-4o with the one-shot prompts to complete the abstract adaptation and reported the scores in BLEU, SARI, BERTScore, LENS, and SALSA.
+In task one (term replacement), we applied fine-tuned ReBERTa-Base models to identify and classify the difficult terms, jargon, and acronyms in the biomedical abstracts and reported the F1 score (Task 1A and 1B).
+In task two (complete abstract adaptation), we leveraged Llamma3.1-70B-Instruct and GPT-4o with the one-shot prompts to complete the abstract adaptation and reported the scores in BLEU, SARI, BERTScore, LENS, and SALSA.
 From the official Evaluation from PLABA-2024 on Task 1A and 1B, our much smaller fine-tuned RoBERTa-Base model ranked 3rd and 2nd respectively on the two sub-tasks, and the 1st on averaged F1 scores across the two tasks from 9 evaluated systems.
-Our LLaMA-3.1-70B-instructed model achieved the highest Completeness score for Task-2.
-We share our source codes, fine-tuned models, and related resources at https://github.com/HECTA-UoM/PLABA-2024
+Our LLaMA-3.1-70B-instructed model achieved the highest Completeness score for Task 2.
+We share our source codes, fine-tuned models, and related resources at https://github.com/HECTA-UoM/PLABA2024
 1 Background
@@
 5 Conclusion and Future Work
-In this report, we present the two systems submitted by the BeeManc team for two different tasks in PLABA-2024.
+In this report, we present the two systems submitted by the MaLei team for two different tasks in PLABA-2024.
 For Task 1, we submitted a classification system based on Roberta for Task 1a and 1b.
@@
 The authors used ChatGPT to refine the paper writing.
+LH and GN were supported by the grant “Integrating hospital outpatient letters into the healthcare data space” (EP/V047949/1; funder: UKRI/EPSRC).
+LH is also grateful to the EU grant 4D Pictures https://4dpicture.eu/.